A new drug on the market is claimed by its manufacturers to reduce overweight women by 4.55 kg per month with a standard deviation of 0.91 kg. Forty women chosen at random have reported losing an average of 4.05 kg within a month.

1
Expert's answer
2022-05-02T03:24:06-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu=4.55"

"H_a:\mu\not=4.55"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.


2. Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is"z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."


3. The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{4.05-4.55}{0.91/\sqrt{40}}\approx-3.475"

4. Since it is observed that "|z| = 3.475 > 1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-3.475)=0.000511," and since "p=0.000511<0.05=\alpha," it is concluded that the null hypothesis is rejected.


5. Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 4.55, at the "\alpha = 0.05" significance level.